### v1

# Towards Next-Generation Urban Decision Support Systems through AI-Powered Generation of Scientific Ontology using Large Language Models - A Case in Optimizing Intermodal Freight Transportation

Abstract
The incorporation of Artificial Intelligence (AI) models into various optimization systems is on the rise. Yet, addressing complex urban and environmental management problems normally requires in-depth domain science and informatics expertise. This expertise is essential for deriving data and simulation-driven for informed decision support. In this context, we investigate the potential of leveraging the pre-trained Large Language Models (LLMs). By adopting ChatGPT API as the reasoning core, we outline an integrated workflow that encompasses natural language processing, methontology-based prompt tuning, and transformers. This workflow automates the creation of scenario-based ontology using existing research articles and technical manuals of urban datasets and simulations. The outcomes of our methodology are knowledge graphs in widely adopted ontology languages (e.g., OWL, RDF, SPARQL). These facilitate the development of urban decision support systems by enhancing the data and metadata modeling, the integration of complex datasets, the coupling of multi-domain simulation models, and the formulation of decision-making metrics and workflow. The feasibility of our methodology is evaluated through a comparative analysis that juxtaposes our AI-generated ontology with the well-known Pizza Ontology employed in tutorials for popular ontology software (e.g., protégé). We close with a real-world case study of optimizing the complex urban system of multi-modal freight transportation by generating anthologies of various domain data and simulations to support informed decision-making.
keywords:
Urban decision support system , Large language models , Ontology , Intermodal freight transportation , Artificial intelligence
1 Introduction
The global urban population has surged from 1.01 billion in 1960 to 4.52 billion in 2022 and is expected to reach 6.9 billion, or approximately 70% of the global population, by 2050 (Hoornweg and Pope, 2017). The accelerated pace of global urbanization underscores the urgent need for effective management and optimization of urban systems. The goals are to promote urban livability through efficient use of resources (Habitat, 2020) and reduce carbon footprint worldwide for more sustainable economic growth under the mounting pressure of climate change (Moran et al., 2018).
Over recent decades, academia, industry, and governmental agencies have significantly invested in developing smart city applications (Angelidou et al., 2018). These efforts aim to foster smart and sustainable cities to enhance urban planning and management and optimize city operations, through the integration of cutting-edge computing and communication technologies (Bibri and Krogstie, 2017). The synergy of these emerging technologies, specifically the Internet of Things (IoT), digital twins, and artificial intelligence (AI), alongside advanced simulations across various urban domains (including transportation, water supply, buildings, and energy systems), fosters a more informed, reliable, and time-sensitive approach to complex urban decision-making (Rane, 2023).
Despite the successes of past smart city applications, challenges persist in optimizing complex urban systems, which involve intricate interactions among various subsystems (McPhearson et al., 2016). A prime example is the intermodal freight transportation system, spanning vast areas and governed by dynamic, multi-scale processes with broad socio-economic and environmental impacts (Heinold and Meisel, 2020).
At the methodological level, the development of effective smart city applications faces formidable challenges, particularly due to the need for interdisciplinary knowledge to create scenario-based ontologies that represent these dynamics accurately (Matei et al., 2021). These ontologies are vital for improving problem understanding, formulating objectives, and designing optimization workflows. From a technical standpoint, integrating diverse urban data and coupling models across domains is crucial for generating both model and data-driven insights, essential for informed decision-making in urban digital twins and decision support systems (Chen et al., 2018). This synthesis supports informed decision-making, making it indispensable in the creation of urban digital twins and decision support systems. Yet, it is often constrained by the laborious nature of the work and the extensive interdisciplinary knowledge required to create and integrate scenario-based ontologies for addressing specific urban challenges.
Figure 1: Conceptual design of an LLM-powered autonomous agent for automated ontology generation.
Recent progress in generative models, notably large language models (LLMs) like GPT-3 (Brown et al., 2020), GPT-4 (Achiam et al., 2023), and Google Gemini (Team et al., 2023) has significantly propelled the research and development of autonomous agents and systems forward. In light of these developments, this paper explores the potential of LLMs to leverage their advanced understanding, reasoning, and generative abilities in human language. Our goal is to embed domain knowledge into scientific ontologies for informed decision support in boosting smart city functionalities. By employing the ChatGPT API as a core reasoning tool, we outline a detailed workflow that incorporates (a) scholarly databases, (b) natural language processing tools, (c) structured prompts from the Methontology workflow, and (d) pre-trained LLM.
We aim to automate the creation of scenario-based ontologies through the analysis of research articles and technical documents on urban data and simulations. Our approach yields knowledge graphs and data dictionaries in standard ontology languages like OWL, RDF, and SPARQL. It enhances aspects of urban decision support systems including data modeling, dataset integration, simulation model linking, and the formulation of decision-making criteria. Further, we evaluate our method’s effectiveness by comparing it with the Pizza Ontology, used in ontology software tutorials such as Protégé. The paper culminates in a case study on optimizing intermodal freight transportation in urban environments, showcasing how our ontologies enable informed decision-making.
2 Literature Review
Ontology and knowledge graphs are foundational elements in the realm of semantic web and knowledge management, offering structured ways to represent knowledge about the world. Defined as “a formal, explicit specification of a shared conceptualisation” by Noy and Hafner (1997), ontology, in a computational sense, is a formal representation of knowledge as a set of concepts within a domain, and the relationships between those concepts (Gruber, 1993). It provides a shared vocabulary for researchers to annotate, share, and reuse their knowledge across different systems and domains (Chen et al., 2018). Knowledge graphs further extend this concept by creating large networks of entities and their interrelations, encapsulated in graph structures, which enable complex queries and analytics, making knowledge accessible and actionable (Ji et al., 2021).
2.1 Ontology for Smart Cities
In smart city applications, ontology and knowledge graphs play pivotal roles in enhancing urban management, decision support, and the optimization of urban systems. They facilitate the integration of diverse data sources, provide a basis for interoperability between different systems and platforms, and support the reasoning processes needed for intelligent decision-making (Chen et al., 2018). By mapping out the complex relationships between various urban elements such as transportation networks, utilities, and services, these tools help in synthesizing data into actionable insights, thereby improving efficiency, sustainability, and the quality of urban life (Bibri and Krogstie, 2017).
Furthermore, ontology and knowledge graphs are crucial in the development of web-based decision support systems (Spoladore and Pessot, 2021), digital twins (Petrova-Antonova and Ilieva, 2021), databases (Balasubramani et al., 2016), AI applications (Bergman et al., 2018), and software dedicated to smart city management (Shi et al., 2023). They provide the underlying architecture for these technologies to dynamically model and simulate the real-world urban environment, enabling cities to anticipate problems, optimize operations, and engage more effectively with citizens (Kuster et al., 2020). Through the use of these semantic technologies, smart city applications, such as digital twin and online decision support systems, can evolve into more adaptive, responsive, and intelligent software applications, harnessing the full potential of the online big urban data to meet the challenges of complex urban system management and optimization (Weil et al., 2023).
2.2 Ontology Design and Development
Methodologies for constructing scientific ontologies employ both detailed and systemic approaches to address the multifaceted nature of ontology development (Cristani and Cuel, 2005). At the micro level, methodologies emphasize the granular aspects of ontology creation, such as the formalization of axioms and design decisions influenced by the chosen representation language (Guizzardi, 2005). Conversely, macro-level approaches provide a comprehensive systems perspective, focusing on the development process through information systems and IT lens (Merali, 2006).
Key methodologies in the field include older frameworks like Methontology and On-To-Knowledge, alongside more contemporary ones that focus on the re-usability and flexibility of ontology, such as NeON.
  • 1.
    Methontology is a rigorous framework developed for the creation of domain ontologies that are independent of applications, facilitating a structured approach to ontology design (Gosal, 2015). This methodology encompasses several key phases, including specification, conceptualization, formalization, implementation, and maintenance, ensuring a comprehensive development process. Originally applied in the late 1990s, Methontology emphasizes systematic steps for the effective organization and representation of knowledge, making it a cornerstone in the field of semantic technologies and knowledge engineering (Fernández-López et al., 1997).
  • 2.
    The On-To-Knowledge methodology (OTKM) represents a structured approach to ontology design, aimed at enhancing knowledge management within organizations. It focuses on converting implicit company knowledge into an explicit, structured form, facilitating better decision-making and process optimization (Sure et al., 2004). This method emphasizes early stakeholder involvement, iterative development, and the use of formal ontologies to ensure that the captured knowledge is both accurate and usable in automated systems, making it a cornerstone for developing semantic applications in various domains.
  • 3.
    The NeON methodology caters to the evolving demands of ontology engineering, supporting the reuse, reengineering, and integration of ontological resources (Gómez-Pérez and Suárez-Figueroa, 2009). By offering a diverse set of strategies and tools, NeON facilitates the development of robust ontologies and ontology networks, addressing complex semantic frameworks (Suárez-Figueroa et al., 2015). This methodology is particularly beneficial for projects requiring collaborative and distributed development, making it a pivotal resource for advancing semantic technologies and their applications in various domains.
The ontology development landscape has evolved towards a collaborative and distributed paradigm, transitioning from the creation of singular, monolithic ontologies to the collaborative construction of ontology networks (Keet, 2004). Consequently, contemporary methodologies adapt to the nuances of collaborative and distributed development, supporting concurrent contributions from domain experts and knowledge engineers across different locales and encouraging a modular approach for enhanced management and scalability (Nguyen, 2011). Each methodology contributes a unique perspective to ontology development, focusing on intricate authoring techniques or overarching system-wide processes. The selection of a methodology can profoundly affect the development trajectory, influenced by variables such as the application domain, team composition, and the intended utility of the ontology.
2.3 Traditional Method for Ontology Implementation
At the implementation level, a variety of software tools for ontology creation, visualization, manipulation, and editing are web-based or desktop-based and are widely utilized by knowledge engineers and domain scientists alike. Among these, Protégé (Noy et al., 2001), OntoSoft (Gil et al., 2016), Hozo (Mizoguchi et al., 2007), and OntoGraf (Falconer, 2010), yEd Graph Editor (Sedlmeier and Gogolla, 2016), Visual Understanding Environment (VUE) (Saigal and Kumar, 2005), and CmapTools (Cañas et al., 2004) have emerged as notable options in recent years. When choosing these tools, ontology developers must consider several critical factors, including the capability for reusing existing ontologies, comprehensive documentation, as well as the ability to export and import data across diverse formats, views, and libraries (Dudáš et al., 2018). These considerations are essential for ensuring the effective application and integration of ontological tools in scientific and engineering workflows.
Implementing ontology with specialized management software is a time-consuming and complex process, largely due to the detailed work required to represent domain-specific knowledge accurately and the expertise needed to effectively utilize these tools (Botzenhardt et al., 2011). The development of ontology is inherently iterative, involving repeated cycles of testing, evaluation, and refinement. Each cycle can reveal new requirements or inconsistencies, necessitating a return to previous stages, which extends the duration of the project (Fernández-López and Gómez-Pérez, 2002). Given these challenges, there is a pressing need to harness emerging AI technologies to automate aspects of ontology implementation, potentially streamlining the process and reducing the manual effort involved.
2.4 Semantic Web Standards
After being implemented, scientific ontologies are typically stored and documented using common semantic web languages that adhere to the semantic web standards, including RDF (Resource Description Framework), OWL (Web Ontology Language), and SPARQL (SPARQL Protocol and RDF Query Language) (Allemang and Hendler, 2011). These standards are crucial for structuring, storing, and querying data effectively. They enable the structured representation and interconnected querying of data across diverse systems, facilitating seamless data integration and accessibility on the Semantic Web.
  • 1.
    RDF serves as a standard model for data interchange on the web, enabling the representation of information in a triple format (subject-predicate-object), facilitating a graph-like structure (Tomaszuk and Hyland-Wood, 2020).
  • 2.
    OWL provides additional vocabulary along with formal semantics designed for processing and integrating information as it enables the creation of more complex ontologies (Cardoso and Pinto, 2015).
  • 3.
    SPARQL, on the other hand, is used to query databases stored in RDF format. It allows for powerful and expressive queries over diverse data sources, making it essential for extracting and manipulating data within and across knowledge graphs (Curé and Blin, 2014).
Together, these semantic web languages form the backbone of semantic web applications, allowing for robust and sophisticated handling of scientific resources, such as the sharing of scientific datasets and modes, across the internet. Many smart city applications depend on semantic web technologies to discover and integrate essential data, thereby facilitating informed urban decision-making and optimizations. Consequently, these technologies play a pivotal role in the development of numerous urban decision support systems. Therefore, it is imperative that future AI-powered techniques for generating ontologies also adhere to established Semantic Web Standards. This alignment will ensure compatibility and enhance the effectiveness of data integration across various platforms and systems.
2.5 AI-powered autonomous system
The swift evolution of AI has catalyzed the emergence of autonomous agents (e.g., industrial robots and autonomous vehicles), also referred to as intelligent agents (Brustoloni, 1991), which are sophisticated computer systems designed to undertake tasks and make decisions autonomously, with little to no human oversight (Wooldridge and Jennings, 1995). These agents have demonstrated significant potential across diverse fields such as robotics, healthcare, transportation, and finance, showcasing benefits like enhanced efficiency, diminished errors, and the capability to manage intricate tasks that might be impractical or exceedingly laborious for humans (Stone et al., 2022). Diverging from automated systems, such as automated irrigation systems and home automation systems, that execute predefined tasks based on specific sets of commands, inputs, rules, and environmental conditions, autonomous agents exhibit higher complexity, better adaptability to their surroundings, and the ability to make informed decisions and content generation from collected data (Albrecht and Stone, 2018).
The capability of most existing rule- or algorithm-based reasoning cores is confined to certain or closed environments, limiting their reasoning capacity for tasks that involve creating new content or computer codes. With the advancement of generative AI models, LLMs like GPT-3 (Brown et al., 2020) and GPT-4 (Sun et al., 2023), has significantly propelled forward the research and development of autonomous agents. These models exhibit a profound comprehension of human natural language, enabling them to adeptly perform tasks across various domains including reasoning, creative writing, code generation, translation, and information retrieval (Li and Ning, 2023).
With recent advancements in prompt engineering and tuning techniques, LLMs are increasingly being integrated as the reasoning core to enable AI-powered assistants or autonomous agents capable of performing specialized tasks, such as generating content within specific domain contexts, automating code generation, and synthesizing data (Shin et al., 2023). This integration highlights the transformative impact of LLMs in enhancing automated processes and driving innovation across various fields, shedding light on an autonomous approach for generating scientific ontology.
2.6 Challenge and Knowledge Gap
Despite the presence of well-established methodologies, ontology management software, and advanced web semantic languages for designing, developing, and implementing scientific ontologies, the creation of efficient and reusable ontologies to support smart city services frequently encounters both methodological and technical challenges. These challenges can impede the effective promotion and integration of such ontologies within the framework of smart city initiatives.
Methodologically, creating scenario-based ontologies tailored to enhance data discovery, integration, and management for specific smart city services, such as parking availability, building control optimization, and traffic volume prediction, is a crucial but repetitive and labor-intensive task in the development of numerous urban decision support tools. This effort becomes increasingly challenging as new data sources and urban simulation scenarios continue to emerge rapidly across expansive urban regions, along with the deployment of new sensors and IoT-connected infrastructure, as well as the increasing data availability from various sources (e.g., crowd-sourcing, autonomous vehicles, and simulation outputs). Moreover, ensuring the reusability and interoperability of newly created scenario-based ontologies remains a significant challenge. For instance, integrating the ontology of a new multimodal freight transportation system with that of an existing urban transportation system can pose complex challenges. The real-time traffic sensor data described in the urban transportation system ontology is also crucial for supporting the operations of the multimodal freight transport and logistics system.
Technically, developing a scenario-based ontology for newly emerging data sources or simulation models often requires time-consuming and labor-intensive efforts. This process demands in-depth domain expertise to extract and interpret essential knowledge from extensive documentation, research articles, and manuals—knowledge that is critical for the specific scope of a smart city application. After extracting the core knowledge, which is subsequently represented as high-level concepts, properties, and relationships, maintaining the semantic consistency of the derived ontology across existing domain ontologies is crucial. This ensures that new data and simulations can be interpreted correctly and uniformly, and readily integrated into the existing software tools. However, this task requires complex and iterative processes, including the standardization of data and metadata, semantic mapping, as well as comprehensive validation and testing.
Given the exceptional capabilities of LLMs in natural language understanding, pattern recognition, reasoning, and self-learning, there is a compelling need to develop an AI-powered autonomous tool for ontology and knowledge graph generation. This tool would leverage LLMs’ capabilities to automate the extraction of knowledge from urban research articles and technical manuals, thereby facilitating the construction of scenario-based ontologies. Such a tool would enhance data-driven research and benefit smart city applications by optimizing complex urban management decisions.
3 Methodology
This section begins with an overview of the background and design requirements for our proposed ontology generation workflow. It then details the development of our methodology, both conceptually and technically. The ontology produced through our approach is validated by comparing it with the well-known Pizza Ontology [^1] Available at: https://drive.google.com/file/d/1PMM-G9KOKrmJwzrqoYnC-VLnAT1fEgZG/view?usp=sharing, which is frequently used in tutorials for popular ontology software.
Figure 2: A subset of sample data from the three datasets combined to demonstrate the origin-destination pairs, as well as different modes of transportation and types of transit across the United States.
3.1 Design Requirement
Our initiative to prototype an LLM-powered autonomous agent for ontology generation is driven by the need to create a large-scale, data-driven digital twin designed to support complex decision-making for optimizing the intermodal transportation system across the contiguous US. Our digital twin, debuted as ARPA-E RECOIL, leverages the latest state-of-the-art intermodal transportation datasets and simulation tools to optimize the freight transport system. This comprehensive approach takes into account supply and demand dynamics, delivery times, and costs, along with potential socioeconomic impacts such as traffic congestion, air pollution, and carbon emissions. Consequently, it necessitates the creation of an ontology behind the proposed urban digital twin to integrate multi-domain datasets and facilitate the coupling simulations across a variety of disciplines. The target users of the proposed autonomous agents include urban researchers, operational scientists, and software developers with domain expertise and either entry-level or intermediate computer science and programming skills. Detailed design requirements of our autonomous agent is as follows:
  • 1.
    Compatibility and Portability: The autonomous agent should be designed as a generalized workflow, implementable in popular programming languages like Python. It should be modular and flexible, enabling deployment and execution on widely-used online platforms such as Jupyter Notebook.
  • 2.
    Automated Knowledge Source Acquisition: The autonomous agent should be integrated with online databases and archives for scientific research literature via APIs. This integration will enable the agent to automatically download relevant research papers and technical manuals based on user-defined keywords related to a specific scope, dataset, or simulation model.
  • 3.
    Autonomous Knowledge Extraction: The autonomous agent should be capable of comprehending human natural language, recognizing domain-specific terminologies, such as identifying concepts as entities and categories as classes, and understanding the relationships between various concepts and classes.
  • 4.
    Automated Ontology Design: The autonomous agent should adhere to the practices and procedures of well-established ontology design methodologies, including Methontology, OTKM, and NeON (see Section 2.2). This adherence will enable it to create robust knowledge representations characterized by high granularity, reusability, flexibility, and semantic consistency.
  • 5.
    Ontology Interoperability: The autonomous agent should possess the capability to export ontologies as knowledge graphs or in widely accepted ontology languages, such as OWL and RDF (see Section 2.4). This functionality will enable the ontologies to be interpreted by popular ontology management software tools, facilitating validation processes and the creation of database schema for data modeling and integration.
  • 6.
    Self-validating: The autonomous agent should have the capability to validate the ontologies it generates by referencing existing domain ontologies or relevant literature. This process should include checking the definitions of concepts, relationships, and semantic mappings to ensure accuracy and consistency.
The autonomous agent, designed and developed based on the aforementioned requirements, will later be employed to create a sample ontology as a proof-of-concept. This will involve integrating emerging datasets and simulation outputs from the intermodal transportation sector, including the Freight Analysis Framework (FAF), Freight Transportation Optimization Tool (FTOT), and Bureau of Transportation Statistics (BTS). This integration aims to facilitate informed and comprehensive decisions to optimize complex intermodal freight transportation systems. More details regarding the data and simulation used in this work are provided in Section 4.1.
3.2 Methodological Foundation
The conceptual design of our autonomous agent is depicted in Figure 3, which comprises four primary modules: (a) Knowledge Source Acquisition, (b) Knowledge Source Pre-processing, (c) LLM-Powered Ontology Generation, and (d) Ontology Implementation. Each component encompasses a series of procedures implemented using Python libraries. Additionally, the entire process is hosted in a Python JupyterLab environment to facilitate human supervision of the workflow step-wise. Detailed workflow and procedures within individual modules are presented through the following subsections.
Figure 3: Overall architecture of the proposed autonomous agent for ontology generation with four major modules
3.2.1 Knowledge Source Acquisition and Pre-Processing
The knowledge source acquisition module automates the discovery, downloading, and archiving of digitized files (e.g., PDFs and XMLs). These files contain essential domain knowledge and are retrieved based on user-defined keywords. These keywords may refer to specific urban or environmental datasets, scientific simulations, and more. The process utilizes SerpApi, a web scraper that identifies and downloads internet resources (e.g., the HTML page of a report or article). Additionally, it employs the Scopus API, an academic database that discovers and downloads research articles using these keywords as search criteria. This API-powered component searches the internet for resources that match the search criteria and downloads them. The resources, including research articles and technical manuals for datasets or simulation models, are saved in common formats such as PDF, HTML, and XML. They are then stored in a designated file directorate, which can be accessed by other components of our autonomous agent.
The knowledge source pre-Processing module is designed to prepare digital files, previously downloaded, into formats suitable for input into the LLM for ontology generation. This module integrates content analysis and data extraction libraries with custom Python scripts to initially cleanse the digital files. This cleansing process removes unnecessary components such as headers, footers, page numbers, and special characters, which could potentially impair the performance of subsequent natural language processing tasks and the LLM’s reasoning capabilities. Subsequently, the content is extracted and transformed into plain text (for text blocks and tables) or images (for figures), depending on their original format. The extracted content is then divided into subsets that comply with the word count limitations of the LLM for each request, specifically using the ChatGPT API in this study.
3.2.2 LLM-powered Ontology Generation
The LLM-powered ontology generation module leverages ChatGPT’s natural language processing (NLP) and reasoning capabilities to automatically generate scenario-based ontologies. This process is facilitated through the use of theory-guided prompts, which are developed based on the core principles of the Methontology approach.
Before initiating the ontology development procedures, plain text and image inputs are processed through two separate workflows to enhance input data quality. Text inputs undergo a series of NLP procedures including tokenization, stop word removal, sentence boundary detection, and text normalization. Concurrently, images are processed through the GPT image recognition engine via API to extract embedded text and summarize the content in text form.
The refined text inputs from both plain texts and images are then fed into the ChatGPT-4 API, which serves as the central reasoning engine, to construct the ontology. The Methontology framework’s procedures have been translated into structured prompts that utilize the flexibility of Python scripting and the ChatGPT API, creating a maintainable workflow. This workflow instructs the LLM on how to create scientific ontologies based on pre-processed input data.
Detailed data flow and its procedure for processing the plain text inputs are described as the following:
  1. 1.
    Stop Word Removal: a technique used in many traditional NLP tasks where common words such as “and”, “the”, “is”, etc., which are deemed to have little meaning, are removed from the text. This is typically done to reduce the dimensionality of the data and focus on more meaningful words for tasks like text classification, keyword extraction, and more, and is conducted outside the ChatGPT API using the NLTK library.
  2. 2.
    An NLP procedure that cleans and standardizes text by converting all characters to lowercase, removing or replacing special characters and punctuation, standardizing dates, numbers, and other non-standard text forms, and managing whitespaces. This process ensures uniformity in the text, minimizing variability that could detract from analysis. Text normalization is integrated into the LLM’s functionality via the ChatGPT API.
  3. 3.
    Sentence Boundary Detection: Following text normalization, this process identifies the start and end points of sentences within the text. Accurate sentence boundary detection is essential for understanding the text’s structure and for facilitating subsequent per-sentence processing. It segments a large text into manageable, analyzable units. This functionality is also integrated into the LLM through the ChatGPT API.
  4. 4.
    Tokenization: Following sentence boundary detection, the text within each sentence is tokenized. This process splits the text into tokens—typically words, phrases, or other meaningful elements. Tokenization transforms the text into discrete units suitable for further processing, such as parsing, part-of-speech tagging, or input to LLMs for reasoning, interpretation, and knowledge extraction. This functionality is also integrated into the LLM through the ChatGPT API.
Detailed data flow and its procedure for processing the image inputs are described as the following:
  1. 1.
    Image Text Extraction: The process uses Optical Character Recognition (OCR) technology to detect and convert text within images into machine-readable characters. In academic figures, this includes extracting annotations, axis labels, data points, and descriptive text, essential for interpreting the graphical content. This process is conducted outside the ChatGPT API using the Pytesseract library.
  2. 2.
    GPT Image Recognition Engine: a component of a larger AI system that utilizes models trained to recognize and interpret content within images. This engine processes images to identify objects, features, and patterns, often using advanced neural networks optimized for image analysis. In the context of research articles, this could involve recognizing graphical data, text within figures, or even the structure of the figures themselves. This functionality is integrated into the LLM through the ChatGPT API.
  3. 3.
    Summary of Images as Texts: The process synthesizes key information from an image into a concise textual summary. It uses AI to interpret recognized objects and text, distilling these elements to highlight the essential insights or data presented. This is especially useful in academic settings for quickly conveying the significance and results depicted in complex figures or diagrams. This functionality is also integrated into the LLM through the ChatGPT API.
This section details the workflow and procedures for leveraging the ChatGPT API’s language understanding and reasoning capabilities, as highlighted by the light-orange border in Figure 3, to develop the ontology. The ontology generation process relies on a series of engineered promotes that are derived and guided by the ontology design procedures from the Methontology framework to guide the LLM in performing specialized tasks. Currently, we do not apply any prompt tuning or fine-tuning technique to our method.
  1. 1.
    Extracting a Glossary of Terms using Named Entity Recognition (NER): This initial step in ontology creation employs NER, a pivotal NLP technique for identifying and categorizing key textual information. The process involves scanning text to detect and label critical domain-specific concepts as entities. These entities form the foundational elements of the ontology, enhancing semantic analysis by clearly outlining the terms and their relationships within the domain.
  2. 2.
    Build Concept Taxonomy: A step focuses on organizing and structuring the identified concepts into a hierarchical framework. Building a concept taxonomy involves arranging concepts (or classes) identified during the conceptualization phase into a taxonomic (hierarchical) structure based on ”is-a” relationships. This hierarchy represents a superclass-subclass relationship between broader and more specific concepts.
  3. 3.
    Define Ad Hoc Relationship: This step is a crucial phase in the ontology development process. It involves identifying and specifying the relationships that exist between the concepts that are specific to the domain being modeled and do not necessarily fit into well-defined hierarchical (is-a) or associative (part-of) relationships.
After creation, the scenario-based ontologies generated by our proposed methods undergo validation against the documentation of existing domain ontologies, which typically describe high-level, generic domain knowledge. This validation is conducted using the reasoning capabilities of LLMs. The purpose of this process is to ensure accurate and relevant identification of domain concepts, as well as logical mapping of relationships between concepts and classes. Additionally, this procedure ensures semantic consistency across the AI-created ontology through a semantic mapping process that leverages the existing domain ontology. After these procedures, the finalized ontology output is then stored in an OWL format, which is interoperable and compatible with popular ontology management software tools and modern semantic web technologies.
3.2.3 The ontology implementation
The OWL ontology produced by our methods is subsequently imported into widely-used ontology management software, Protégé as part of an existing project. This process supports a range of use cases that enhance the development of decision support systems or digital twins by providing direct or indirect pathways to generate essential software components. Examples of these components include data models and queries, knowledge graphs, and ontologies stored in other popular Semantic Web formats (Dong et al., 2007).
A specific use case involves converting the LLM-generated OWL ontology into SPARQL (SPARQL Protocol and RDF Query Language) using Protégé. SPARQL is primarily used to query and manipulate data stored in the Resource Description Framework (RDF) format (Kumar and Kumar, 2013). This capability enables decision support applications to more effectively discover, retrieve, and integrate data resources from RDF databases. In this setting, the ontology can be used to facilitate data management from relational database (Lubbad, 2018), enabling an data provision pipeline to support data-driven insights during the development of urban decision support system.
3.3 Qualitative Result Evaluation
(a) Outer Level
(b) Inner Level
Figure 4: We show a human-created ontology, which defines a subclass-of relationship. This relationship, expressed as “A is-a B,” signifies that A represents a more specific or specialized category than B. In essence, this arrangement delineates the hierarchical structuring of knowledge.
We evaluate the feasibility of our methodology through a comparative analysis, juxtaposing our AI-generated ontology with the widely recognized Pizza Ontology. Class materials [^2] Available at: http://www.meteck.org/swdsont.html and the work discussed by Rector et al. (2004), depict the challenges faced in the domain of web ontology language description and generation logic. These challenges often can lead to conceptual misunderstandings and underscore the necessity for expert domain knowledge to effectively structure, comprehend, and ultimately construct knowledge representations.
In the realm of the use of an automated ontology creation process, we qualitatively verified its proper generation, which contrasts with the Pizza representation, to ensure the LLM-generated ontology have valid . We focus on two sections of this ontology; under Figure 4, we present the human-created ontology, which shows an “is-a” or “subclass-of” relationship. This subtype or subclass, indicates a specialization-generalization relationship.
It’s crucial to highlight the current empirical evidence that underscores the robust capabilities of applying methodologies akin to those outlined in Fernández-López et al. (1997), now seamlessly executed through Large Language Models (LLMs). The ontology illustrated in Figure 4 amalgamates inputs from undisclosed sources, revealing the logical interconnections among the presented classes and subclasses. Conversely, the data utilized for the automated pizza ontology stems from a distinct singular source, presented in a plain format.
The modular components outlined in the subsections of the ontology pizza source document delineate a systematic breakdown of information, subsequently processed through the proposed framework and LLM. Figures 5(b) and 5(a) delineate pertinent sections and mirror the same logical components and properties showcased in the diagram.
(a) Outer Level
(b) Inner Level
Figure 5: An unprocessed automated ontology per section fully generated via the LLM. This ontology depicts vegetable toppings embedded as individuals in the LLM generated model. The inclusion of individual instances composed by artichokes, mushrooms, onion tomatoes resembles the human made ontology.
Under 5(b) we represented a narrow display of detail level represented via a pizza class and corresponding sub classes that carry styles, characteristic tool and components. In addition, label under ’Pizza’ a deeper level can be seen. In 5(a) This more general level of understanding about the context of pizza
4 Case Study: Ontology in Intermodal Freight Transportation
In this section, we present a compelling use case illustrating the application of our methodology in tackling a real-world decision optimization challenge. Our focus lies in the intricate realm of intermodal freight transportation across the United States, where we harness the power of integrating multiple datasets and simulations. The primary objective of the study is to optimize this complex system with the goals of enhancing fuel efficiency and, at the same time, reducing Greenhouse Gas (GHG) emissions and total operation costs. This endeavor stands as a demonstration to the capability of our methodology in addressing multifaceted, real-world challenges with tangible environmental and economic implications in urban systems.
4.1 Data Sources and Simulation Models
Our methodology integrates three primary data sources alongside external data from academic sources, enriching the knowledge and analysis of intermodal transportation networks as represented under Figure 3. The Freight Analysis Framework (FAF) (U.S. Department of Transportation, 2024) and its accompanying manual serve as foundational resources, identifying crucial freight transportation hubs across diverse regions. These hubs provide a framework for mapping significant geographic and logistic parameters that define major nodes within transportation networks. Similarly, the FTOT (U.S. Department of Energy, 2024b) and its manual are instrumental in permitting the routing engine to perform calculations under the criteria of the shortest path; this transportation network is represented via shapefiles. The generated ontology will permit LLMs to perform calculations and handle services such as an endpoint REST API to compute unique routes for various nodes, encompassing road, rail, and water modes, while also delineating the properties and relationships within these networks. Manuals ensure a comprehensive understanding of the methodologies and algorithms applied in these tools, enhancing the accuracy of simulations and optimizations in our model.
The Alternative Fueling Stations (AFS-BTS) dataset (U.S. Department of Energy, 2024a) averages 1,521 locations per state and encompasses a total of 80,601 locations across the US providing valuable insights into alternative fueling stations. Of particular interest are the 69,314 electric charging stations, highlighting electric vehicles as the most prevalent renewable fuel type. The dataset also includes information on fuel types including Electric, E85, Liquefied Petroleum Gas, Bio-Diesel, Compressed Natural Gas, Renewable Diesel, Liquefied Natural Gas, and Hydrogen. This dataset facilitates detailed modeling of CO2 emissions, a critical aspect of our optimization and emission reduction strategies. We incorporate a fourth data source. Academic contributions, such as methodological approaches for calculating equivalent CO2 emissions, among other sources, enhance our understanding and integration of emissions considerations within the LLM framework, further refining our strategies for optimization and sustainability in intermodal freight transportation.
In addition to the aforementioned primary data and simulations encompassing operational costs, fuel consumption, and GHG emissions across various modes of freight transportation in the US, our approach integrates a comprehensive array of open GIS datasets. These datasets delineate critical fuel and transportation infrastructure locations, leveraging resources such as the Homeland Infrastructure Foundation-Level Data (HIFLD) for fuel station and supply locations, and the OpenStreetMap network for vital highways and rail systems. Furthermore, to capture the dynamic physical processes inherent in intermodal freight transportation, we incorporate real-time traffic congestion and accident information. This invaluable data is sourced from real-time traffic sensor feeds accessible via web APIs. Noteworthy examples include the Radar Detection System operated by the Tennessee Department of Transportation, along with the WAZE API.
4.2 Ontology-driven Decision Optimization
The methodology, empowered by LLM, initially utilized for creating the pizza ontology, is now repurposed to generate a scenario-based knowledge representation. This representation integrates various intermodal freight transportation data and simulation outputs, alongside other urban datasets. The resulting real-time decision support workflow targets the optimization of urban logistic systems across extensive geographic areas, demonstrating the versatility and efficacy of the LLM-powered approach, as depicted in Figure 6.
Figure 6: Decision support workflow that is driven by the scenario-based ontology created using the LLM-powered autonomous workflow.
The case study presented in this section utilizes a subset of FAF, FTOT, and AFS BTS data to establish a use case for identifying the optimal combination of shipping modes and routes from Nashville to New Orleans. Traditional manual approaches to decision optimization often demand extensive domain expertise regarding the dynamics of inter-modal freight transportation events, as well as data- and simulation-driven insights for quantifying various decision support criteria, such as operational costs and emissions.
Employing our prototyping methodology for creating scenario-based ontology, we have successfully developed a digital representation of knowledge sourced from the technical manuals of major intermodal freight transportation datasets and simulations. This digital knowledge representation is exported into an OWL format, furnishing an ontology-driven pipeline to delineate the decision support workflow and to structure data and simulation outputs within the domain context. Our ontology consists of three major components, as the following:
  • a)
    Graph Representation is a graph data model that captures the topological properties (connectivity) of the freight transportation network. It is derived by the LLM using information from the FAF dataset manual. The model represents various modes and routes of freight transportation between key origin-destination pairs, which are represented by US cities that are major logistic and transportation hubs. This graph is primarily used to navigate between any user-defined origin-destination pairs.
  • b)
    Ontology Representation delineates the major entities, along with their relationships and properties, within the intermodal freight transportation network as represented by the graph model. This ontology is derived from the information outlined in the manuals of both the FAF and FTOT datasets. It provides a foundational data structure that maps data variables—used to characterize the physical dynamics and performance of intermodal transportation processes—to their corresponding graph entities.
  • c)
    Data Sources comprise all the necessary multi-domain datasets required to provide data- and simulation-driven insights, complete with essential decision variables. These insights aid in making informed decisions aimed at optimizing operational costs and delivery times, while also reducing fuel consumption and GHG emission. These data sources are organized through a scenario-based ontology and are mapped to corresponding network entities (nodes and edges). Subsequently, they are ingested into a relational database, utilizing a data schema derived from the ontology to organize the data to enable rapid data retrieval for decision support purposes.
Leveraging these ontological components, this pipeline facilitates the data warehousing, management, and decision workflow formulation during the development of a data- and simulation-driven decision support system for intermodal freight transportation optimization. The pipeline entails four steps:
  • a)
    The “Network Traverse” refers to the systematic process of navigating through a transportation network to identify all possible combinations of routes and modes between an origin-destination pair. This process involves assessing various transportation modes—such as road, rail, and river—and their interconnections, along with transit locations represented as intermediate nodes within the FTOT network. Each mode and route combination is characterized by a series of nodes that signify the origin, destination, and intermodal transits, connected by edges that represent the transitions between modes. This comprehensive approach ensures a thorough exploration of the network, facilitating optimal route and mode selection.
  • b)
    The “Decision Support Variable Definition” is a critical process that identifies key variables essential for characterizing the physical processes involved in intermodal transportation. These variables are vital for determining the most efficient, cost-effective, or fastest combinations of routes and modes available. During the ontology creation stage with LLM, structured prompts are developed. These prompts guide the ChatGPT API in extracting relevant variables from the technical manuals of the FAF and FTOT datasets, focusing on the efficiency and sustainability aspects of the freight transportation system. This structured approach ensures that all critical factors are considered in decision-making processes, enhancing the system’s overall performance and sustainability.
  • c)
    The “Data and Model Integration” is facilitated by encoding the Uniform Resource Identifier (URI) of specific data or simulations within a relational database, accessible via an API or hosted as a web service, into the ontology as properties of the graph entities. This method ensures that the freight transportation data and simulations are readily available during the network traversal. As the network is traversed to identify an individual combination of route and mode, decision variables at each graph entity are aggregated into decision metrics for the entire route combination, such as total greenhouse gas (GHG) emissions and total operational costs. This integrated setup enhances the efficiency of accessing and utilizing data, thereby optimizing the decision-making process in freight transportation management. The aggregated decision metrics for each route and mode combination are consolidated into a relational lookup table, which serves as a foundation for further decision analysis.
  • d)
    The “Optimal Mode and Route Identification” process utilizes optimization algorithms on a lookup table filled with decision metrics to systematically evaluate and select the most effective transportation options. This evaluation considers multiple, potentially conflicting criteria such as GHG emissions, operational costs, fuel consumption, and shipment time. By integrating both network optimization and Multi-Criteria Decision Analysis, this approach ensures a comprehensive and reliable determination of the optimal route and mode, balancing various transportation needs and environmental impacts effectively.
Our methodology capitalizes on the capabilities of LLMs to automate the digitization of domain knowledge encapsulated in text and figures within technical manuals of urban datasets, simulations, and research articles. This knowledge is converted into interoperable ontology formats, which underpin the development of data models. These models structure essential data and simulation outputs to support decision-making processes. The lookup table, generated in real-time from the ontology-driven data model, plays a pivotal role. Together, they form crucial components for the development of urban decision support systems or digital twin applications, enhancing the efficiency and effectiveness of urban planning and management.
5 Limitation and Future Work
Throughout the course of development and evaluation, we have identified two major limitations associated with our current AI-powered methodology for generating ontologies. These limitations, along with their potential solutions through future work, are outlined as follows:
  1. 1.
    Without precisely engineered prompts, the ChatGPT API may exhibit sensitivity, leading to subtle variability and inconsistent results across different iterations. The robustness of the model can be enhanced by employing more advanced techniques, such as fine-tuning and prompt tuning. These methods adapt pre-trained LLMs to generate ontologies using specific datasets or under particular conditions, thereby minimizing potential variability in the results from different model runs.
  2. 2.
    There is a limitation on the maximum number of tokens that the ChatGPT API can process per request. This constraint can be circumvented by partitioning, optimizing, and refining the input data, combined with advancements in LLM research.
We aim to address these limitations through our future work.
6 Conclusion
In this study, we introduced an AI-powered methodology designed to automate the generation of scenario-based scientific ontologies by harnessing the text understanding and reasoning capabilities of LLMs. This methodology was implemented into a prototype autonomous agent using Python NLP libraries and the ChatGPT API. Our primary goal is to support the development of urban decision support systems by automating time-consuming and labor-intensive tasks, which typically require substantial domain knowledge. Our approach facilitates the integration of data resources and simulations to enhance informed decision-making.
We tested our methodology by generating the well-known Pizza Ontology and compared the AI-generated outputs with the expert-created Pizza Ontology, validating the reliability and feasibility of our approach. Through a case study, we demonstrated how our methodology could be applied to real-world urban data, such as the FAF and FTOT datasets, to aid in the development of an ontology-driven urban decision support system. Although still in its early stages, our approach showcases the potential and feasibility of using an autonomous agent to generate ontologies that support scientific software and database development. This initiative sheds light on the next generation of decision support systems that leverage large-scale urban data to optimize complex urban systems.
With significant promise for achieving autonomous decision support systems, we believe our methodology can be further extended and applied to the creation of various domain ontologies, thereby lowering technical barriers and enhancing the capability to solve complex urban system challenges. Additionally, our methods can be leveraged to automate bibliometric analyses, focusing on the extraction of key domain concepts. Future work of this study includes applying more recent model tuning techniques, such as the prompt tuning and fine tuning, to further instruct LLM on generating more specialized ontology, as well as software modules and codes that serve as the critical components of a urban decision support system.
7 Acknowledgements
This work was supported in part by the U.S. Department of Energy’s Advanced Research Projects Agency-Energy (ARPA-E) under the project (#DE-AR0001780) titled “A Cognitive Freight Transportation Digital Twin for Resiliency and Emission Control Through Optimizing Intermodal Logistics” (RECOIL).